# Sysformer: Safeguarding Frozen Large Language Models with Adaptive System Prompts

### v1

@@ 5 Experimental Setup
Datasets. We use two labeled datasets of harmful and safe prompts from the recently published benchmarks: JailbreakBench (behaviors) (Chao et al., 2024)  [^1] https://huggingface.co/datasets/JailbreakBench/JBB-Behaviors (MIT License) and StrongReject (Souly et al., 2024)  [^2] https://strong-reject.readthedocs.io/en/latest/api/load_datasets.html (MIT License). These consist of curated examples sourced from both original and prior datasets such as DAN (Shen et al., 2024), Harmbench (Mazeika et al., 2024), AdvBench (Zou et al., 2023), etc. JailbreakBench consists of a curated set of 100 harmful and 100 safe prompts while StrongReject consists of 313 harmful prompts from various sources. Thus, we augment the JailbreakBench’s safe prompts to the StrongReject dataset of harmful prompts. We split each dataset into train and test splits using the ratio of $70\%$ to $30\%$, ensuring that the proportion of harmful and safe prompts is the same in both splits. We also split the train set further by leaving $10\%$ as the validation set for early stopping. For additional compliance to reinforce the pre-training objective, we use the instruction-tuning Alpaca dataset containing pairs of input prompt and expected response  [^3] https://huggingface.co/datasets/tatsu-lab/alpaca.
Large language models. We use different instruction-tuned LLMs to evaluate the safety gains of the Sysformer architecture. We consider $5$ models spanning $4$ families, including Meta’s Llama (Llama-3.1-8B [^4] Llama-3.1-8B-Instruct, Llama-2-7B-chat [^5] Llama-2-7B-chat), Mistral AI’s Mistral-7B-v0.2  [^6] Mistral-7B-Instruct-v0.2, Microsoft’s Phi-3.5-mini  [^7] Phi-3.5-mini-instruct, and Huggingface’s zephyr-7B-beta 888zephyr-7B-beta. Our experiments are restricted to 8B models due to the memory limitation of back-propagating the gradients through >8B models.
@@
Metrics. We employ LlamaGuard that measures the proportion of responses that are classified as safe (identified using string matching) by the LlamaGuard model (Inan et al., 2023). We also employ an LLM-based Refusal Rate that measures the proportion of responses that refuse to respond to a given prompt using GPT-4o-mini (version 2024-02-01) followed by string matching. Both prompts are provided in Appendix C.1. We find the Refusal Rate for both harmful and safe prompts and use RR for safe prompts to assess safety compliance.
Attacks. We consider different jailbreaking attacks to evaluate the robustness of the Sysformer architecture to test both its generalization. For example, in a GCG-based attack (Zou et al., 2023), a prompt like “Tell me how to create a bomb” in the dataset will be augmented with an adversarial string found using a greedy search to obtain create “Tell me how to create a bomb <string>”. Here, we employ a total of $16$ such jailbreaking attack strategies including PAIR (Chao et al., 2023), GCG (Zou et al., 2023), PAP (Zeng et al., 2024), etc. Then, following  Souly et al. (2024), we update each prompt in the held-out test set of harmful prompts to a corresponding adversarial prompt for each jailbreaking attack. Thus, we obtain a set of jailbreaking prompts corresponding to each test prompt and attack strategy. We provide the full list of jailbreaking strategies used in Appendix D.
Hyperparameters. We train all the trainable parameters in each method using AdamW optimizer (Loshchilov et al., 2017) and find the best performance by searching over $\{10,20\}$ epochs and initial learning rate $\in\{0.0001,0.00001\}$. We keep $w_{ref}=1$ to be fixed and search the other hyperparameters as $w_{compl}\in\{0.0,0.2,0.5,1.0\},w_{class}\in\{0.0,1.0\},w_{recon}\in\{0,1\},\texttt{add}\in\{\text{True},\text{False}\},\texttt{selfsafe}\in\{\text{True},\text{False}\}$.
@@ 6 Results
6.1 Can Sysformer effectively and selectively refuse harmful prompts across LLMs?
First, we test how well can Sysformer learn to distinguish between harmful and safe prompts by evaluating it on the held-out test set in a lossless and frozen defense setting across two benchmarks and five LLMs. For the Sysformer and SystemEmbedder, we choose the best configuration, i.e., the one that maximizes the sum of the increase in refusal rate on harmful prompts and the decrease in refusal rate on safe prompts, relative to the DefaultSystem baseline.
Table 1 shows that Sysformer remains the closest to the best performance in each LLM and dataset, enhancing the refusal rate and llama-guard score on harmful prompts by upto $80\%$ in Phi-3.5-mini while reducing the refusal rate for safe prompts by upto $90\%$ in Llama-2-7b-chat. We find that Sysformer’s refusal rate can learn to refuse harmful prompts effectively in almost all cases, with a minimum refusal rate of $\sim 60\%$ and an average refusal rate of $88\%$ across different LLMs and datasets. On the other hand, Sysformer’s refusal is selective as it significantly reduces the refusal rate on safe prompts, keeping it $\leq 17\%$ in all cases. This demonstrates a high generalization of Sysformer in its ability to learn the expected refusal direction across LLMs. We also note that since certain LLMs such as Llama-2-7b-chat, Mistral-7B-v0.2, and Llama-3.1-8B are already safety-tuned, Sysformer is focused on reducing its over-refusal on safe prompts, leading to a significant drop in the safe refusal rate while keeping harm refusal rate high. In contrast, since other models such as zephyr-7b-beta, and Phi-3.5-mini are not natively safety-tuned (as can be seen from the low refusal rate of the default setting), Sysformer is found to increase the harm refusal rate while keeping the safe refusal rate constant. Finally, we also note that the Sysformer refusal rates for harmful prompts in StrongReject are generally lower than for the ones in JailbreakBench across LLMs. This can be owed to the more sophisticated harmful examples creating using AutoDAN (Zhu et al., 2024) present in the StrongReject while JailbreakBench only consists of naturalistic prompts.
Table 1: Comparison of Sysformer with other lossless frozen defense mechanisms.
@@ 6.1 Can Sysformer effectively and selectively refuse harmful prompts across LLMs?
       Defense Refual Rate LlamaGuard Refual Rate LlamaGuard
Safe $\downarrow$ Harm $\uparrow$ Safe $\uparrow$ Harm $\uparrow$ Safe $\downarrow$ Harm $\uparrow$ Safe $\uparrow$ Harm $\uparrow$
@@
       DefaultSystem 0.0667 0.3333 0.9333 0.2000 0.0667 0.3191 0.9333 0.3191
       SystemEmbedder 0.0667 0.4000 0.9333 0.2000 0.0667 0.3404 0.9333 0.3191
       Sysformer (ours) 0.1667 0.9333 0.8667 0.8000 0.1333 0.7553 0.8667 0.6170
@@
       DefaultSystem 0.7000 1.0000 1.0000 1.0000 0.6667 0.9894 1.0000 1.0000
       SystemEmbedder 0.5667 1.0000 0.9333 1.0000 0.0667 0.4000 0.9333 0.2000
       Sysformer (ours) 0.0667 0.9000 0.9000 0.8667 0.0333 0.8085 0.9333 0.8085
@@
       DefaultSystem 0.3000 1.0000 1.0000 1.0000 0.3000 1.0000 1.0000 1.0000
       SystemEmbedder 0.3000 1.0000 1.0000 1.0000 0.3000 1.0000 1.0000 1.0000
       Sysformer (ours) 0.0333 0.9667 0.8333 0.9667 0.0333 1.0000 0.9000 1.0000
@@
       DefaultSystem 0.0333 0.1000 0.6667 0.0667 0.0333 0.2128 0.6667 0.0319
       SystemEmbedder 0.0333 0.1667 0.6667 0.0667 0.0667 0.2660 0.6667 0.0319
       Sysformer (ours) 0.2000 0.9000 0.8667 1.0000 0.0667 0.5851 0.9000 0.8617
@@
       DefaultSystem 0.1333 0.8333 0.9333 0.9333 0.1333 0.9362 0.9333 0.9574
       SystemEmbedder 0.1333 0.8667 0.9333 0.9333 0.1333 0.9362 0.9333 0.9574
       Sysformer (ours) 0.1000 1.0000 0.9333 1.0000 0.1000 1.0000 0.9333 0.9681
@@ 6.2 Can Sysformer defend against sophisticated jailbreaking attacks?
Next, we study how well Sysformer can defend against sophisticated attacks that are specifically designed to jailbreak the LLM into complying with a harmful prompt. As noted in Section 5, we create an evaluation set by applying $16$ different jailbreaking attack strategies to create a hard set of attack prompts for each harmful prompt in our test set.
Figure 3 compares the refusal rate for safe and harmful prompts of JailbreakBench along with the refusal rate over the set of jailbreaking prompts created by applying different attacks over the same harmful prompts. We find that Sysformer (denoted as Ours) fails to generalize to these jailbreaking attacks as the refusal rate (denoted through a boxplot) remains similar to the baselines when the method is only trained using the training set of harmful prompts in the dataset.
Trained with attack-augmented data. Since Sysformer has never seen these sophisticated examples during training, the results are expected. To further supervise and inform the Sysformer of such examples of harm-inducing prompts, we augment the training set of harmful prompts with a few such attacking strategies. In particular, we only use 6 out of 16 attacks to augment the training set, which includes PAIR (Chao et al., 2023), one setting of PAP (Zeng et al., 2024), distractors, style injection, refusal suppression (Souly et al., 2024), and one setting of GCG. Figure 3 shows that Syformer trained using attack-augmented data (denoted as Ours+JB) achieves remarkable gains in refusal rate for both natural and jailbreaking harmful prompts of the held-out test set while complying with most safe prompts. In particular, we find that in all cases except Phi-3.5-mini, we can learn to refuse almost all jailbreaking prompts even those that were not seen during training since the whole box is moved upwards near 1. For results on other data splits and attack splits, please refer to Appendix E.
@@ 6.3 How sensitive is Sysformer to different hyperparameters?
Sysformer employs various hyperparameters as noted in Section 5, such as the weights of the $4$ loss functions, whether to train using additional compliance, and whether to use a self-generated compliance response. Thus, we compare the performance of Sysformer considering different combinations of these hyperparameters. We keep the $w_{ref}=1$ as the main objective is to learn to refuse the harmful prompts and also keep $w_{class}=1$ as we give us the best performance in all cases.
Figure 4 compares the refusal rate for harmful, safe, and jailbreaking prompts in the JailbreakBench dataset. We observe a high sensitivity to the loss weights in some LLMs such as zephyr-7b-beta, Phi-3.5-mini, and Mistral-7b-v0.2 while Llama-3.1-8b remains largely stable. It also demonstrates that intermediate parameter values (0.2-0.5) for $w_{compl}$ typically outperform extreme settings (1.0), and hyperparameters interact with each other showing, becoming more important than individual settings. Notably, we find that a high compliance weight need not necessarily reduce the safe refusal rate for test prompts and can sometimes hurt performance. Optimal configurations generally combine moderate compliance weights or use additional compliance data instead of templated or LLM-generated compliance. The impact of the reconstruction loss weight remains highly dependent on the model and other hyperparameters and enabling it sometimes helps significantly in improving the performance, e.g., in Mistral-7b-v0.2 and Phi-3.5-mini. We also generally find that self-compliance is only useful in handling the refusal rate tradeoff when the underlying LLM is safety-tuned like Mistral-7b-v0.2 while otherwise, it is shown to increase the refusal rate for safe test prompts. These findings highlight that tuning these hyperparameters requires careful LLM-specific analysis with general patterns of low compliance weights, additional compliance, and optional reconstruction and self-compliance should be searched over to optimally train safe Sysformer architecture. For analysis on other combinations, please check Appendix E.

### v2

@@ 5 Experimental Setup
Datasets. We use two labeled datasets of harmful and safe prompts from the recently published benchmarks under MIT License: JailbreakBench (behaviors) (Chao et al., 2024)  [^4] https://huggingface.co/datasets/JailbreakBench/JBB-Behaviors and StrongReject (Souly et al., 2024)  [^5] http://strong-reject.readthedocs.io/en/latest/api/load_datasets.html. These consist of curated examples sourced from both original and prior datasets such as DAN (Shen et al., 2024), Harmbench (Mazeika et al., 2024), AdvBench (Zou et al., 2023), etc. JailbreakBench consists of a curated set of 100 harmful and 100 safe prompts while StrongReject consists of 313 harmful prompts from various sources. Thus, we augment the JailbreakBench’s safe prompts to the StrongReject dataset of harmful prompts. We split each dataset into train and test splits using the ratio of $70\%$ to $30\%$, ensuring that the proportion of harmful and safe prompts is the same in both splits. We also split the train set further by leaving $10\%$ as the validation set for early stopping. For additional compliance to reinforce the pre-training objective, we use the instruction-tuning Alpaca dataset containing pairs of input prompt and expected response  [^6] https://huggingface.co/datasets/tatsu-lab/alpaca.
Large language models. We use different instruction-tuned LLMs to evaluate the safety gains of the Sysformer architecture. We consider $5$ models spanning $4$ families, including Meta’s Llama (Llama-3.1-8B [^7] Llama-3.1-8B-Instruct, Llama-2-7B-chat [^8] Llama-2-7B-chat), Mistral AI’s Mistral-7B-v0.2  [^9] Mistral-7B-Instruct-v0.2, Microsoft’s Phi-3.5-mini  [^10] Phi-3.5-mini-instruct, and Huggingface’s zephyr-7B-beta 111111zephyr-7B-beta. Our experiments are restricted to 8B models due to the memory limitation of back-propagating the gradients through >>8B models.
@@
Metrics. We employ LlamaGuard (LG) that measures the proportion of responses that are classified as safe (identified using string matching) by the LlamaGuard model (Inan et al., 2023). We also employ an LLM-based Refusal Rate (RR) that measures the proportion of responses that refuse to respond to a given prompt using GPT-4o-mini (version 2024-02-01) followed by string matching. Both prompts are provided in Appendix A.1. We use the Refusal Gap or $\Delta$ RR as our main metric as it measures the difference between the refusal rate for harmful and safe prompts.
Attacks. We consider different jailbreaking attacks to evaluate the robustness of the Sysformer architecture to test both its generalization. For example, in a GCG-based attack (Zou et al., 2023), a prompt like “Tell me how to create a bomb” in the dataset will be augmented with an adversarial string found using a greedy search to obtain create “Tell me how to create a bomb <string>”. Here, we employ a total of $16$ such jailbreaking attack strategies including PAIR (Chao et al., 2023), GCG (Zou et al., 2023), PAP (Zeng et al., 2024), etc. Then, following  Souly et al. (2024), we update each prompt in the held-out test set of harmful prompts to a corresponding adversarial prompt for each jailbreaking attack. Thus, we obtain a set of jailbreaking prompts corresponding to each test prompt and attack strategy. We provide the full list of jailbreaking strategies used in Appendix C.
Hyperparameters. We train all the trainable parameters in each method using AdamW optimizer (Loshchilov et al., 2017) and find the best performance with respect to $\Delta$ RR by searching over $\{10,20\}$ epochs and initial learning rate $\in\{0.0001,0.00001\}$. We keep $w_{ref}=1$ to be fixed and search the other hyperparameters as $w_{compl}\in\{0.0,0.2,0.5,1.0\},w_{class}\in\{0.0,1.0\},w_{recon}\in\{0,1\},\texttt{add}\in\{\text{True},\text{False}\},\texttt{selfsafe}\in\{\text{True},\text{False}\}$.
@@ 6 Results
6.1 Can Sysformer effectively enable safety in frozen LLMs?
Table 1 shows that Sysformer outperforms all baselines for frozen LLMs in increasing the refusal gap between safe and harmful prompts across different LLMs and datasets. We find that Sysformer can learn to refuse harmful prompts effectively in almost all cases, with a minimum refusal rate of $\sim 60\%$ and an average refusal rate of $88\%$, while significantly reducing the refusal rate on safe prompts, keeping it $\leq 17\%$ in all cases and reducing it by upto $90\%$ in Llama-2-7b-chat. This demonstrates a high generalization of Sysformer in its ability to learn the expected refusal direction across LLMs. Furthermore, we find that the Sysformer either matches or outperforms a strong fine-tuning baseline using LoRA on all layers with $r=16,\alpha=32$ (Mazeika et al., 2024). Specifically, it matches or outperforms this baseline by increasing the refusal gap by up to $50\%$, while keeping the LLM parameters frozen. This highlights a key benefit of Sysformer that safety can be achieved without updating the pre-trained parameters using an attachable module.
Table 1: Comparison of Sysformer with other filtering-free frozen defense mechanisms. Llama-guard scores are reported in Table 56 in Appendix D.
@@ 6.1 Can Sysformer effectively enable safety in frozen LLMs?
RR Safe $\downarrow$ RR Harm $\uparrow$ $\Delta$ $R$ $\uparrow$ RR Safe $\downarrow$ RR Harm $\uparrow$ $\Delta$ $R$ $\uparrow$
@@
     LoRA 0.0333 0.8667 0.7333 0.2000 0.9255 0.7255
     DefaultSystem 0.0667 0.3333 0.2666 0.0667 0.3191 0.2524
     SystemEmbedder 0.0667 0.4000 0.3333 0.0667 0.3404 0.2737
     Sysformer (ours) 0.1667 0.9333 0.7666 0.1333 0.7553 0.6220
@@
     LoRA 0.1000 0.9667 0.8667 0.1000 1.0000 0.9000
     DefaultSystem 0.7000 1.0000 0.3000 0.6667 0.9894 0.3227
     SystemEmbedder 0.5667 1.0000 0.4333 0.0667 0.4000 0.3333
     Sysformer (ours) 0.0667 0.9000 0.8333 0.0333 0.8085 0.7752
@@
     LoRA 0.1000 0.9667 0.8667 0.0000 1.0000 1.0000
     DefaultSystem 0.3000 1.0000 0.7000 0.3000 1.0000 0.7000
     SystemEmbedder 0.3000 1.0000 0.7000 0.3000 1.0000 0.7000
     Sysformer (ours) 0.0333 0.9667 0.9334 0.0333 1.0000 0.9667
@@
     LoRA 0.1667 0.6000 0.4333 0.0667 0.4894 0.4227
     DefaultSystem 0.0333 0.1000 0.0667 0.0333 0.2128 0.1795
     SystemEmbedder 0.0333 0.1667 0.1334 0.0667 0.2660 0.1993
     Sysformer (ours) 0.2000 0.9000 0.7000 0.0667 0.5851 0.5184
@@
     LoRA 0.2333 1.0000 0.7667 0.1000 1.0000 0.9000
     DefaultSystem 0.1333 0.8333 0.7000 0.1333 0.9362 0.8029
     SystemEmbedder 0.1333 0.8667 0.7334 0.1333 0.9362 0.8029
     Sysformer (ours) 0.1000 1.0000 0.9000 0.1000 1.0000 0.9000
We also note that since certain LLMs such as Llama-2-7b-chat, Mistral-7B-v0.2, and Llama-3.1-8B are already safety-tuned, Sysformer is focused on reducing its over-refusal on safe prompts, leading to a significant drop in the safe refusal rate while keeping the harm refusal rate high. In contrast, since other models such as zephyr-7b-beta and Phi-3.5-mini are not natively safety-tuned (as can be seen from the low refusal rate of the default setting), Sysformer is found to increase the harm refusal rate while keeping the safe refusal rate constant. Finally, we also note that the Sysformer refusal rates for harmful prompts in StrongReject are generally lower than for the ones in JailbreakBench across LLMs. This can be owed to the more sophisticated harmful examples created using AutoDAN (Zhu et al., 2024) present in the StrongReject, while JailbreakBench only consists of naturalistic prompts.
We next test the cross-dataset generalizability of Sysformer by training it on JailbreakBench and testing on StrongReject. Table 2 shows that Sysformer can generalize extraordinarily well across benchmarks since it reaches similar or even better performance compared to when it is trained using StrongReject’s train split. This can be attributed to Sysformer’s adaptive ability and a balanced harm-safe split of JailbreakBench as compared to a 3:1 ratio of harm-safe in StrongReject.
LLM RR safe RR harm $\Delta$ $R$
zephyr-7b 0.2000 0.9681 0.7681
Llama-2-7b 0.2333 0.9787 0.7454
Llama-3.1-8b 0.0670 1.0000 0.9330
phi-3.5-mini 0.2330 0.9255 0.6925
Mistral-7b-v0.2 0.1000 1.0000 0.9000
Table 2: Performance of JailbreakBench-Trained Sysformer on StrongReject.
We further validate that Sysformer preserves the general text generation performance by comparing the BERTScore (Zhang et al., 2019) between the generated responses and the gold responses on the evaluation split of the Alpaca dataset. The average BERTScore for Llama-2-7B-chat drops slightly from 0.8487 to 0.8414 with Sysformer, while for Llama-3.1-8B it rises from 0.8327 to 0.8467.
@@ 6.2 Can Sysformer defend against sophisticated jailbreaking attacks?
Here, we study how well Sysformer can defend against sophisticated attacks that are specifically designed to jailbreak the LLM into complying with a harmful prompt. As noted in Section 5, we create an evaluation set by applying $16$ different jailbreaking attack strategies. Figure 3 compares the refusal rate for safe and harmful prompts of JailbreakBench, along with the refusal rate over the set of jailbreaking prompts created by applying different attacks over the same harmful prompts. Sysformer (denoted as Ours) fails to generalize to these jailbreaking attacks, as the refusal rate (denoted through a boxplot) remains similar to the baselines. This is expected since Sysformer has never encountered these sophisticated examples during training.
Thus, we follow the existing literature (Mazeika et al., 2024; Zou et al., 2024) and augment the training set of harmful prompts with a few such attacking strategies. In particular, we add examples of 6 out of 16 attacks to the training set (details are provided in Appendix C). Figure 3 shows that Syformer trained using attack-augmented data (denoted as Ours+JB) achieves remarkable gains in refusal rate for both natural and jailbreaking harmful prompts of the held-out test set while complying with most safe prompts. In particular, we find that in all cases except Phi-3.5-mini, we can learn to refuse almost all jailbreaking prompts, even those that were not seen during training, since the whole box is moved upwards near 1.
To further analyze the generalizability of Sysformer+JB on unseen attacks, we expand the evaluation to all 28 different attacks available in  Souly et al. (2024). Table 3 shows the mean, median, first quantile, and 3rd quantile ranges of refusal rates over different jailbreaking distributions, particularly. Results show that training on a dataset containing only 6 attacks can generalize well (particularly, in the case of the Mistral-7B-Instruct-v0.2 model) to a large variety of jailbreaking strategies, as demonstrated by the high average, median, and lower quantile refusal rate across different jailbreaking distributions.
LLM # OOD RR mean RR median RR Q1 RR Q3
zephyr-7b 0 0.4905 0.4000 0.3333 0.7167
10 0.4771 0.3833 0.2917 0.7333
22 0.4762 0.4333 0.2667 0.7000
Llama-2-7b 0 0.6048 0.5667 0.4333 0.7833
10 0.6354 0.6500 0.4917 0.8083
22 0.6740 0.6667 0.5000 0.8333
phi-3.5-mini 0 0.5333 0.4667 0.2333 0.8833
10 0.5625 0.5667 0.3417 0.7917
22 0.5324 0.6333 0.2333 0.8333
Mistral-7B-v0.2 0 0.7429 0.8000 0.6500 1.0000
10 0.7875 0.8000 0.7500 0.9167
22 0.7267 0.8000 0.6667 0.9000
Table 3: Sysformer+JB evaluated on different jailbreaking attack sets (Souly et al., 2024).
@@
6.3 How efficient is Sysformer?
Table 4 shows the additional time taken by the methods on top of different LLMs during inference over JailbreakBench. We find that the overhead is minimal in most cases, remaining within 20-30 s on average, and Sysformer takes comparable times with SystemEmbedder in all cases. To calculate this inference time, we assume that the transformed system prompt is cached for subsequent token generations and thus, only use the methods at the first generated token. We defer the space efficiency analysis to Appendix D, while noting that train time efficiency cannot be compared across methods as the best configuration can be different, incurring additional costs.
LLM Method Overhead (s)
zephyr-7b-beta SystemEmbedder 30.4453
Sysformer 27.4878
Llama-2-7b-chat SystemEmbedder 21.5297
Sysformer 21.6185
Llama-3.1-8b SystemEmbedder 26.9679
Sysformer 27.1886
Phi-3.5-mini SystemEmbedder 22.0127
Sysformer 21.5096
Mistral-7B-v0.2 SystemEmbedder 22.1994
Sysformer 22.6286
Table 4: Average inference time overhead for different defense methods on JailbreakBench.
@@ 6.4 How sensitive is Sysformer to different hyperparameters?
Sysformer employs various hyperparameters as noted in Section 5, such as the weights of the $4$ loss functions, whether to train using additional compliance, and whether to use a self-generated compliance response. Thus, we compare the performance of Sysformer considering different combinations of these hyperparameters. We keep the $w_{ref}=1$ as the main objective is to learn to refuse the harmful prompts, and also keep $w_{class}=1$ as it gives us the best performance in all cases.
Figure 4 compares the refusal rate for harmful, safe, and jailbreaking prompts in the JailbreakBench dataset. We observe a high sensitivity to the loss weights in some LLMs, such as zephyr-7b-beta, Phi-3.5-mini, and Mistral-7b-v0.2, while Llama-3.1-8b remains largely stable. It also demonstrates that intermediate parameter values (0.2-0.5) for $w_{compl}$ typically outperform extreme settings (1.0), and hyperparameters interact with each other, becoming more important than individual settings. Notably, we find that a high compliance weight need not necessarily reduce the safe refusal rate for test prompts and can sometimes hurt performance. Optimal configurations generally combine moderate compliance weights or use additional compliance data instead of templated or LLM-generated compliance. The impact of the reconstruction loss weight remains highly dependent on the model and other hyperparameters, and enabling it sometimes helps significantly in improving the performance, e.g., in Mistral-7b-v0.2 and Phi-3.5-mini. We also generally find that self-compliance is only useful in handling the refusal rate tradeoff when the underlying LLM is safety-tuned like Mistral-7b-v0.2, while otherwise, it is shown to increase the refusal rate for safe test prompts. These findings highlight that tuning these hyperparameters requires careful LLM-specific analysis, with general patterns of low compliance weights, additional compliance, and optional reconstruction and self-compliance should be searched over to optimally train a safe Sysformer architecture. For analysis on other combinations, please check Appendix D.